OpenStudy (anonymous):

Which of the following expressions could give the total amount of a $45 restaurant bill with a 15% tip added on? Will medal and fan.

OpenStudy (babynini):

which are the expressions? :)

OpenStudy (anonymous):

a. (0.15)($45) b. (1.15)($45) c. 1 + (0.15)($45) d. $45 + 0.15

OpenStudy (babynini):

do you have any ideas?

OpenStudy (anonymous):

I think it is a. or d. but I have a feeling that its wrong.

OpenStudy (babynini):

okies. hmm, maybe if we try finding out what 15% of 45 is. do you know how to find that?

OpenStudy (anonymous):

No, not really.

OpenStudy (babynini):

ok :) this is sort of a cheater way but.here we go to find out 1% you divide 45 by 100 does it make sense why?

OpenStudy (anonymous):

No, not really. Why?

OpenStudy (babynini):

because 100% is the full amount so 45 = 100% in this case.

OpenStudy (anonymous):

Oh, yeah now I understand. Then what do I do?

OpenStudy (babynini):

Sorry pc got super stuck D:

OpenStudy (babynini):

so you have 45/100 = ?

OpenStudy (anonymous):

45/100 = 0.45. Oh! We have to add 0.45 with 0.15! 0.45+ 0.15 = 0.60 or 0.6! Thank You!

OpenStudy (babynini):

did you get it? haha which letter answer did you get?

OpenStudy (babynini):

well we want to make sure our answer doesn't give us just the tip money but the tip money + 45$

OpenStudy (anonymous):

I got d.

OpenStudy (anonymous):

Is it right?

OpenStudy (babynini):

not quite

OpenStudy (babynini):

ok so 45/100 = will give you the 1% what does 45/100 = ?

OpenStudy (anonymous):

45/100 = 0.45

OpenStudy (babynini):

good. Now, since we want 15% and not just 1% you need to do 0.45 x 15 = ?

OpenStudy (anonymous):

0.45 x 15 = 6.75?

OpenStudy (babynini):

Yes!! so $6.75 is how much the tip is

OpenStudy (babynini):

Now, the question asks for the equation that gives the $45 + tip or $45 + 6.75 so that = ?

OpenStudy (anonymous):

So we have to do $45 + $6.75?

OpenStudy (babynini):

yes! :) what does that equal?

OpenStudy (anonymous):

Oh! $51.75! Right?! Thank you!

OpenStudy (babynini):

Yes. So now find the equation that gives you that answer

OpenStudy (anonymous):

Thank you a million times!

OpenStudy (babynini):

What was your answer now? :)

OpenStudy (anonymous):

I got c. Is it right?